# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ Acknowledgements.
- Halder et al. (2025) Sajal Halder, Muhammad Ejaz Ahmed, and Seyit Camtepe. 2025. FuncVul: An Effective Function Level Vulnerability Detection Model using LLM and Code Chunk. arXiv:2506.19453
@@
- Jiang et al. (2025) Xuefeng Jiang, Lvhua Wu, Sheng Sun, Jia Li, Jingjing Xue, Yuwei Wang, Tingting Wu, and Min Liu. 2025. Investigating Large Language Models for Code Vulnerability Detection: An Experimental Study. arXiv:2412.18260
@@
- Khan et al. (2025) Arshiya Khan, Guannan Liu, and Xing Gao. 2025. Code Vulnerability Repair with Large Language Model Using Context-Aware Prompt Tuning. In Security and Privacy Workshops (SPW). IEEE, San Francisco, CA, USA, 283–287. doi:10.1109/SPW67851.2025.00040
@@
- Li et al. (2025b) Jiayuan Li, Lei Cui, Jie Zhang, Haiqiang Fei, Yu Chen, and Hongsong Zhu. 2025b. Steering Large Language Models for Vulnerability Detection. In International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, Hyderabad, India, 1–5. doi:10.1109/ICASSP49660.2025.10887736
@@
- Li et al. (2025c) Xinghang Li, Jingzhe Ding, Chao Peng, Bing Zhao, Xiang Gao, Hongwan Gao, and Xinchen Gu. 2025c. SafeGenBench: A Benchmark Framework for Security Vulnerability Detection in LLM-Generated Code. arXiv:2506.05692
@@
- Li et al. (2025e) Yue Li, Xiao Li, Hao Wu, Minghui Xu, Yue Zhang, Xiuzhen Cheng, Fengyuan Xu, and Sheng Zhong. 2025e. Everything You Wanted to Know About LLM-based Vulnerability Detection But Were Afraid to Ask. arXiv:2504.13474
- Li et al. (2025f) Yikun Li, Ting Zhang, Ratnadira Widyasari, Yan Naing Tun, Huu Hung Nguyen, Tan Bui, Ivana Clairine Irsan, Yiran Cheng, Xiang Lan, Han Wei Ang, Frank Liauw, Martin Weyssow, Hong Jin Kang, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025f. CleanVul: Automatic Function-Level Vulnerability Detection in Code Commits Using LLM Heuristics. arXiv:2411.17274
- Li et al. (2025d) Ziyang Li, Saikat Dutta, and Mayur Naik. 2025d. IRIS: LLM-Assisted Static Analysis for Detecting Security Vulnerabilities. arXiv:2405.17238
@@
- Lingxiang et al. (2025) Wang Lingxiang, Quanzhi Fu, Wenjia Song, Gelei Deng, Yi Liu, Dan Williams, and Ying Zhang. 2025. SAVANT: Vulnerability Detection in Application Dependencies through Semantic-Guided Reachability Analysis. arXiv:2506.17798
- Liu et al. (2025b) Ethan TS. Liu, Austin Wang, Spencer Mateega, Carlos Georgescu, and Danny Tang. 2025b. VADER: A Human-Evaluated Benchmark for Vulnerability Assessment, Detection, Explanation, and Remediation. arXiv:2505.19395
- Liu et al. (2024b) Fang Liu, Yang Liu, Lin Shi, Houkun Huang, Ruifeng Wang, Zhen Yang, Li Zhang, Zhongqi Li, and Yuchi Ma. 2024b. Exploring and Evaluating Hallucinations in LLM-Powered Code Generation. arXiv:2404.00971
@@
- Liu et al. (2023b) Jiaxing Liu, Chaofeng Sha, and Xin Peng. 2023b. An Empirical Study of Parameter-Efficient Fine-Tuning Methods for Pre-Trained Code Models. In International Conference on Automated Software Engineering (ASE). IEEE/ACM, Kirchberg, Luxembourg, 397–408. doi:10.1109/ASE56229.2023.00125
- Liu et al. (2025c) Ruitong Liu, Yanbin Wang, Haitao Xu, Jianguo Sun, Fan Zhang, Peiyue Li, and Zhenhao Guo. 2025c. Vul-LMGNNs: Fusing Language Models and Online-distilled Graph Neural Networks for Code Vulnerability Detection. Information Fusion 115 (2025), 102748. doi:10.1016/j.inffus.2024.102748
- Liu et al. (2024c) Shigang Liu, Bushra Sabir, Seung Ick Jang, Yuval Kansal, Yansong Gao, Kristen Moore, Alsharif Abuadbba, and Surya Nepal. 2024c. From Solitary Directives to Interactive Encouragement! LLM Secure Code Generation by Natural Language Prompting. arXiv:2410.14321
- Liu et al. (2024a) Yu Liu, Lang Gao, Mingxin Yang, Yu Xie, Ping Chen, Xiaojin Zhang, and Wei Chen. 2024a. VulDetectBench: Evaluating the Deep Capability of Vulnerability Detection with Large Language Models. arXiv:2406.07595
@@
- Liu et al. (2023a) Zhihong Liu, Qing Liao, Wenchao Gu, and Cuiyun Gao. 2023a. Software Vulnerability Detection with GPT and In-Context Learning. In International Conference on Data Science in Cyberspace (DSC). IEEE, Hefei, China, 229–236. doi:10.1109/DSC59305.2023.00041
@@
- Manuel et al. (2024) Dylan Manuel, Nafis Tanveer Islam, Joseph Khoury, Ana Nunez, Elias Bou-Harb, and Peyman Najafirad. 2024. Enhancing Reverse Engineering: Investigating and Benchmarking Large Language Models for Vulnerability Analysis in Decompiled Binaries. arXiv:2411.04981
- Mao et al. (2025) Qiheng Mao, Zhenhao Li, Xing Hu, Kui Liu, Xin Xia, and Jianling Sun. 2025. Towards Explainable Vulnerability Detection with Large Language Models. arXiv:2406.09701
@@
- Mou et al. (2025) Yutao Mou, Xiao Deng, Yuxiao Luo, Shikun Zhang, and Wei Ye. 2025. Can You Really Trust Code Copilots? Evaluating Large Language Models from a Code Security Perspective. arXiv:2505.10494
@@
- Nguyen et al. (2024) Van Nguyen, Surya Nepal, Tingmin Wu, Xingliang Yuan, and Carsten Rudolph. 2024. SAFE: Advancing Large Language Models in Leveraging Semantic and Syntactic Relationships for Software Vulnerability Detection. arXiv:2409.00882
@@
- Ni et al. (2024a) Chao Ni, Liyu Shen, Xiaodan Xu, Xin Yin, and Shaohua Wang. 2024a. Learning-based Models for Vulnerability Detection: An Extensive Study. arXiv:2408.07526
- Ni et al. (2024b) Chao Ni, Liyu Shen, Xiaohu Yang, Yan Zhu, and Shaohua Wang. 2024b. MegaVul: A C/C++ Vulnerability Dataset with Comprehensive Code Representations. In International Conference on Mining Software Repositories (MSR). ACM, Lisbon, Portugal, 738–742. doi:10.1145/3643991.3644886
@@
- Nunez et al. (2024) Ana Nunez, Nafis Tanveer Islam, Sumit Kumar Jha, and Peyman Najafirad. 2024. AutoSafeCoder: A Multi-Agent Framework for Securing LLM Code Generation through Static Analysis and Fuzz Testing. arXiv:2409.10737
@@
- Pan et al. (2021) Cong Pan, Minyan Lu, and Biao Xu. 2021. An Empirical Study on Software Defect Prediction Using CodeBERT Model. Applied Sciences 11, 11 (2021), 4793.
@@
- Qi et al. (2024) Weiliang Qi, Jiahao Cao, Darsh Poddar, Sophia Li, and Xinda Wang. 2024. Enhancing Pre-Trained Language Models for Vulnerability Detection via Semantic-Preserving Data Augmentation. arXiv:2410.00249
@@
- Safdar et al. (2023) Rijha Safdar, M. Umer Ashfaq, and Danyail Mateen. 2023. Deep Learning-Based Framework for Automated Vulnerability Detection in Android Applications. In International Bhurban Conference on Applied Sciences and Technology (IBCAST). IEEE, Bhurban, Murree, Pakistan, 1–5. doi:10.1109/IBCAST59916.2023.10713017
- Sahar et al. (2024) Sadia Sahar, Muhammad Younas, Muhammad Murad Khan, and Muhammad Umer Sarwar. 2024. DP-CCL: A Supervised Contrastive Learning Approach Using CodeBERT Model in Software Defect Prediction. IEEE Access 12 (2024), 22582–22594. doi:10.1109/ACCESS.2024.3362896
@@
- Seo et al. (2025) Minjae Seo, Wonwoo Choi, Myoungsung You, and Seungwon Shin. 2025. AutoPatch: Multi-Agent Framework for Patching Real-World CVE Vulnerabilities. arXiv:2505.04195
@@
- Shashwat et al. (2024) Kumar Shashwat, Francis Hahn, Xinming Ou, Dmitry Goldgof, Lawrence Hall, Jay Ligatti, S. Raj Rajgopalan, and Armin Ziaie Tabari. 2024. A Preliminary Study on Using Large Language Models in Software Pentesting. arXiv:2401.17459
- Sheng et al. (2025) Ze Sheng, Zhicheng Chen, Shuning Gu, Heqing Huang, Guofei Gu, and Jeff Huang. 2025. LLMs in Software Security: A Survey of Vulnerability Detection Techniques and Insights. arXiv:2502.07049
@@
- Song et al. (2023) Wei Song, Lu Gan, and Tie Bao. 2023. Software Defect Prediction via Code Language Models. In International Conference on Communication Technology and Information Technology (ICCTIT). IEEE, Xi’an, China, 97–102. doi:10.1109/ICCTIT60726.2023.10435711
@@
- Sun et al. (2025) Yuqiang Sun, Daoyuan Wu, Yue Xue, Han Liu, Wei Ma, Lyuye Zhang, Yang Liu, and Yingjiu Li. 2025. LLM4Vuln: A Unified Evaluation Framework for Decoupling and Enhancing LLMs’ Vulnerability Reasoning. arXiv:2401.16185
@@
- Szabó and Bilicki (2023) Zoltán Szabó and Vilmos Bilicki. 2023. A New Approach to Web Application Security: Utilizing GPT Language Models for Source Code Inspection. Future Internet 15, 10 (2023), 326.
@@
- Uddin et al. (2022) Md Nasir Uddin, Bixin Li, Zafar Ali, Pavlos Kefalas, Inayat Khan, and Islam Zada. 2022. Software Defect Prediction Employing BiLSTM and BERT-based Semantic Feature. Soft Computing 26, 16 (2022), 7877–7891.
@@
- Wang et al. (2023e) Chaozheng Wang, Yuanhang Yang, Cuiyun Gao, Yun Peng, Hongyu Zhang, and Michael R. Lyu. 2023e. Prompt Tuning in Code Intelligence: An Experimental Evaluation. IEEE Transactions on Software Engineering (TSE) 49, 11 (2023), 4869–4885. doi:10.1109/TSE.2023.3313881
@@
- Wang et al. (2024e) Rongcun Wang, Senlei Xu, Yuan Tian, Xingyu Ji, Xiaobing Sun, and Shujuang Jiang. 2024e. SCL-CVD: Supervised Contrastive Learning for Code Vulnerability Detection via GraphCodeBERT. Computers & Security (COSE) 145 (2024), 103994. doi:10.1016/j.cose.2024.103994
- Wang et al. (2025c) Weizhou Wang, Eric Liu, Xiangyu Guo, Xiao Hu, Ilya Grishchenko, and David Lie. 2025c. ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data. arXiv:2408.16028
@@
- Wang et al. (2024d) Xuanye Wang, Lu Lu, Zhanyu Yang, Qingyan Tian, and Haisha Lin. 2024d. Parameter-efficient Multi-classification Software Defect Detection Method based on Pre-trained LLMs. International Journal of Computational Intelligence Systems (IJCIS) 17, 1 (2024), 152. doi:10.1007/s44196-024-00551-3

### v2

@@ Acknowledgements.
- Hairan and Şahman (2025) Borhanullah Hairan and Mehmet Akif Şahman. 2025. Compressing Large Language Models for SQL Injection Detection: A Case Study on Deep Seek-Coder and Meta-llama-3-70b-instruct. In Conference of Open Innovations Association (FRUCT). FRUCT Oy, Helsinki, Finland, 90–96. doi:10.23919/FRUCT67853.2025.11239157
- Halder et al. (2025) Sajal Halder, Muhammad Ejaz Ahmed, and Seyit Camtepe. 2025. FuncVul: An Effective Function Level Vulnerability Detection Model using LLM and Code Chunk. In European Symposium on Research in Computer Security (ESORICS). Springer, Toulouse, France, 166–185. doi:10.1007/978-3-032-07884-1_9
- Han and Zheng (2025) Yaolin Han and Wei Zheng. 2025. An Automatic Classification Model for Long Code Vulnerabilities Based on the Teacher-Student Framework. In International Conference on Software Quality, Reliability, and Security (QRS) Companion. IEEE, Hangzhou, China, 472–482. doi:10.1109/QRS-C65679.2025.00065
@@
- Hannan et al. (2025) Md Abdul Hannan, Ronghao Ni, Chi Zhang, Limin Jia, Ravi Mangal, and Corina S. Pasareanu. 2025. On Selecting Few-Shot Examples for LLM-based Code Vulnerability Detection. arXiv:2510.27675
- Haque et al. (2025) Radowanul Haque, Aftab Ali, Sally McClean, and Naveed Khan. 2025. A Zero-Shot Framework for Cross-Project Vulnerability Detection in Source Code. Empirical Software Engineering (EMSE) 31, 1 (2025), 3. doi:10.1007/s10664-025-10749-4
@@
- Jararweh et al. (2025) Ala Jararweh, Michael Adams, Avinash Sahu, Abdullah Mueen, and Afsah Anwar. 2025. LLaVul: A Multimodal LLM for Interpretable Vulnerability Reasoning about Source Code. In Intelligent Cybersecurity Conference (ICSC). IEEE, Tampa, FL, USA, 232–241. doi:10.1109/ICSC65596.2025.11140501
@@
- Jiang et al. (2025b) Xuefeng Jiang, Lvhua Wu, Sheng Sun, Jia Li, Jingjing Xue, Yuwei Wang, Tingting Wu, and Min Liu. 2025b. Investigating Large Language Models for Code Vulnerability Detection: An Experimental Study. arXiv:2412.18260
- Jiang et al. (2022) Xuxiang Jiang, Yinhao Xiao, Jun Wang, and Wei Zhang. 2022. Multi-view Pre-trained Model for Code Vulnerability Identification. In International Conference on Wireless Algorithms, Systems, and Applications (WASA). Springer, Dalian, China, 127–135. doi:10.1007/978-3-031-19211-1_11
- Jiang et al. (2025a) Yuan Jiang, Zhichen Qu, Christoph Treude, Xiaohong Su, and Tiantian Wang. 2025a. Enhancing Fine-Grained Vulnerability Detection With Reinforcement Learning. IEEE Transactions on Software Engineering (TSE) 51, 10 (2025), 2900–2920. doi:10.1109/TSE.2025.3603400
@@
- Jiao et al. (2025) Yutong Jiao, Jiaxuan Han, and Cheng Huang. 2025. DeepVulHunter: Enhancing the Code Vulnerability Detection Capability of LLMs through Multi-Round Analysis. Journal of Intelligent Information Systems (JIIS) 63, 6 (2025), 2237–2264. doi:10.1007/s10844-025-00982-0
@@
- Lekssays et al. (2025) Ahmed Lekssays, Hamza Mouhcine, Khang Tran, Ting Yu, and Issa Khalil. 2025. LLMxCPG: Context-Aware Vulnerability Detection Through Code Property Graph-Guided Large Language Models. In 34th USENIX Security Symposium (USENIX Security). USENIX, Seattle, WA, USA, 489–507.
@@
- Li et al. (2025c) Jiayuan Li, Lei Cui, Wenyan Yu, Haiqiang Fei, Feng Cheng, and Hongsong Zhu. 2025c. VulnTeam: A Team Collaboration Framework for LLM-based Vulnerability Detection. In International Joint Conference on Neural Networks (IJCNN). IEEE, Rome, Italy, 1–8. doi:10.1109/IJCNN64981.2025.11229292
- Li et al. (2025d) Jiayuan Li, Lei Cui, Jie Zhang, Haiqiang Fei, Yu Chen, and Hongsong Zhu. 2025d. Steering Large Language Models for Vulnerability Detection. In International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, Hyderabad, India, 1–5. doi:10.1109/ICASSP49660.2025.10887736
- Li et al. (2025e) Jiayuan Li, Lei Cui, Sen Zhao, Yun Yang, Lun Li, and Hongsong Zhu. 2025e. CLeVeR: Multi-modal Contrastive Learning for Vulnerability Code Representation. In Findings of the Association for Computational Linguistics (ACL). ACL, Vienna, Austria, 7940–7951. doi:10.18653/v1/2025.findings-acl.414
@@
- Li et al. (2025b) Yikun Li, Ngoc Tan Bui, Ting Zhang, Martin Weyssow, Chengran Yang, Xin Zhou, Jinfeng Jiang, Junkai Chen, Huihui Huang, Huu Hung Nguyen, Chiok Yew Ho, Jie Tan, Ruiyin Li, Yide Yin, Han Wei Ang, Frank Liauw, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025b. Out of Distribution, Out of Luck: How Well Can LLMs Trained on Vulnerability Datasets Detect Top 25 CWE Weaknesses? arXiv:2507.21817
- Li et al. (2025h) Youpeng Li, Kartik Joshi, Xinda Wang, and Eric Wong. 2025h. MAVUL: Multi-Agent Vulnerability Detection via Contextual Reasoning and Interactive Refinement. arXiv:2510.00317
- Li et al. (2025i) Yue Li, Xiao Li, Hao Wu, Minghui Xu, Yue Zhang, Xiuzhen Cheng, Fengyuan Xu, and Sheng Zhong. 2025i. Everything You Wanted to Know About LLM-based Vulnerability Detection But Were Afraid to Ask. arXiv:2504.13474
- Li et al. (2025j) Youpeng Li, Weiliang Qi, Xuyu Wang, Fuxun Yu, and Xinda Wang. 2025j. Revisiting Pre-trained Language Models for Vulnerability Detection. arXiv:2507.16887
- Li et al. (2025k) Yikun Li, Ting Zhang, Ratnadira Widyasari, Yan Naing Tun, Huu Hung Nguyen, Tan Bui, Ivana Clairine Irsan, Yiran Cheng, Xiang Lan, Han Wei Ang, Frank Liauw, Martin Weyssow, Hong Jin Kang, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025k. CleanVul: Automatic Function-Level Vulnerability Detection in Code Commits Using LLM Heuristics. arXiv:2411.17274
- Li et al. (2025f) Ziyang Li, Saikat Dutta, and Mayur Naik. 2025f. IRIS: LLM-Assisted Static Analysis for Detecting Security Vulnerabilities. In International Conference on Learning Representations (ICLR). ICLR, Singapore, Singapore.
- Li et al. (2025g) Zhihao Li, Zimo Ji, Tao Zheng, Hao Ren, and Xiao Lan. 2025g. CryptoScope: Utilizing Large Language Models for Automated Cryptographic Logic Vulnerability Detection. arXiv:2508.11599
@@
- Liang et al. (2024) Chen Liang, Qiang Wei, Zirui Jiang, Yisen Wang, and Jiang Du. 2024. A Source Code Vulnerability Detection Method Based on Adaptive Graph Neural Networks. In IEEE/ACM International Conference on Automated Software Engineering (ASE) Workshops. IEEE, Sacramento, CA, USA, 187–196.
@@
- Liu et al. (2024a) Chongyang Liu, Xiang Chen, Xiangwei Li, and Yinxing Xue. 2024a. Making Vulnerability Prediction more Practical: Prediction, Categorization, and Localization. Information and Software Technology (IST) 171 (2024), 107458. doi:10.1016/j.infsof.2024.107458
- Liu et al. (2024c) Fang Liu, Yang Liu, Lin Shi, Houkun Huang, Ruifeng Wang, Zhen Yang, Li Zhang, Zhongqi Li, and Yuchi Ma. 2024c. Exploring and Evaluating Hallucinations in LLM-Powered Code Generation. arXiv:2404.00971
@@
- Liu et al. (2025b) Ruitong Liu, Yanbin Wang, Haitao Xu, Jianguo Sun, Fan Zhang, Peiyue Li, and Zhenhao Guo. 2025b. Vul-LMGNNs: Fusing Language Models and Online-distilled Graph Neural Networks for Code Vulnerability Detection. Information Fusion 115 (2025), 102748. doi:10.1016/j.inffus.2024.102748
- Liu et al. (2026) Xiangyue Liu, Xinwei Liu, Lili Bo, Xiaoxue Wu, Yun Yang, Xiaobing Sun, and Feng Zhou. 2026. PIONEER: Improving the Robustness of Student Models when Compressing Pre-Trained Models of Code. Automated Software Engineering (ASE) 33, 1 (2026), 1–30. doi:10.1007/s10515-025-00560-2
- Liu et al. (2024b) Yu Liu, Lang Gao, Mingxin Yang, Yu Xie, Ping Chen, Xiaojin Zhang, and Wei Chen. 2024b. VulDetectBench: Evaluating the Deep Capability of Vulnerability Detection with Large Language Models. arXiv:2406.07595
@@
- Liu et al. (2023) Zhihong Liu, Qing Liao, Wenchao Gu, and Cuiyun Gao. 2023. Software Vulnerability Detection with GPT and In-Context Learning. In International Conference on Data Science in Cyberspace (DSC). IEEE, Hefei, China, 229–236. doi:10.1109/DSC59305.2023.00041
@@
- Maligazhdarova et al. (2025) Nargiz Maligazhdarova, Avinash B. M., Assel Mukasheva, Didar Yedilkhan, Aidos Askhatuly, and A. Berdyshev. 2025. A Comparative Study of Machine Learning and Large Language Models for SQL and NoSQL Injection Vulnerability Detection. In IEEE International Conference on Smart Information Systems and Technologies (SIST). IEEE, Astana, Kazakhstan, 1–7. doi:10.1109/SIST61657.2025.11139190
@@
- Mao et al. (2025) Qiheng Mao, Zhenhao Li, Xing Hu, Kui Liu, Xin Xia, and Jianling Sun. 2025. Towards Explainable Vulnerability Detection With Large Language Models. IEEE Transactions on Software Engineering (TSE) 51, 10 (2025), 2957–2971. doi:10.1109/TSE.2025.3605442
@@
- Mhatre et al. (2025) Akshay Mhatre, Noujoud Nader, Patrick Diehl, and Deepti Gupta. 2025. LLM-GUARD: Large Language Model-Based Detection and Repair of Bugs and Security Vulnerabilities in C++ and Python. arXiv:2508.16419
@@
- Mock et al. (2025) Moritz Mock, Thomas Forrer, and Barbara Russo. 2025. Cross-Domain Evaluation of Transformer-Based Vulnerability Detection on Open and Industry Data. In International Conference on Product-Focused Software Process Improvement (PROFES). Springer, Salerno, Italy, 36–52. doi:10.1007/978-3-032-12092-2_3
@@
- Ngan et al. (2025) Nguyen Dac Thien Ngan, Nghi Hoang Khoa, Van-Hau Pham, and Phan The Duy. 2025. Multimodal Fusion for Vulnerability Detection: Integrating Sequence and Graph-Based Analysis with LLM Augmentation. In International Conference on Multimedia Analysis and Pattern Recognition (MAPR). IEEE, Nha Trang, Vietnam, 1–6. doi:10.1109/MAPR67746.2025.11133833
- Nguyen et al. (2025a) Hong Quy Nguyen, Thong Hoang, Hoa Khanh Dam, Guoxin Su, Zhenchang Xing, Qinghua Lu, and Jiamou Sun. 2025a. Human-Understandable Explanation for Software Vulnerability Prediction. Journal of Systems and Software (JSS) 228 (2025), 112455. doi:10.1016/j.jss.2025.112455
- Nguyen et al. (2025c) Van Nguyen, Surya Nepal, Xingliang Yuan, Tingmin Wu, Fengchao Chen, and Carsten Rudolph. 2025c. MulVuln: Enhancing Pre-trained LMs with Shared and Language-Specific Knowledge for Multilingual Vulnerability Detection. arXiv:2510.04397
- Nguyen et al. (2025b) Van Nguyen, Surya Nepal, Xingliang Yuan, Tingmin Wu, and Carsten Rudolph. 2025b. SAFE: A Novel Approach For Software Vulnerability Detection from Enhancing The Capability of Large Language Models. In ACM Asia Conference on Computer and Communications Security (ASIACCS). ACM, Ha Noi, Vietnam, 392–406. doi:10.1145/3708821.3736208
@@
- Ni et al. (2024) Chao Ni, Liyu Shen, Xiaohu Yang, Yan Zhu, and Shaohua Wang. 2024. MegaVul: A C/C++ Vulnerability Dataset with Comprehensive Code Representations. In International Conference on Mining Software Repositories (MSR). ACM, Lisbon, Portugal, 738–742. doi:10.1145/3643991.3644886
- Ni et al. (2025) Chao Ni, Xin Yin, Liyu Shen, and Shaohua Wang. 2025. Learning-based Models for Vulnerability Detection: An Extensive Study. Empirical Software Engineering (EMSE) 31, 18 (2025), 38. doi:10.1007/s10664-025-10734-x
@@
- Ouchebara and Dupont (2025) Dyna Soumhane Ouchebara and Stéphane Dupont. 2025. Llama-Based Source Code Vulnerability Detection: Prompt Engineering vs Fine Tuning. In European Symposium on Research in Computer Security (ESORICS). Springer, Toulouse, France, 289–308. doi:10.1007/978-3-032-07884-1_15
@@
- Pan et al. (2024) Shirui Pan, Linhao Luo, Yufei Wang, Chen Chen, Jiapu Wang, and Xindong Wu. 2024. Unifying Large Language Models and Knowledge Graphs: A Roadmap. IEEE Transactions on Knowledge and Data Engineering (TKDE) 36, 7 (2024), 3580–3599. doi:10.1109/TKDE.2024.3352100
@@
- Peng et al. (2025) Yun Peng, Kisub Kim, Linghan Meng, and Kui Liu. 2025. iCodeReviewer: Improving Secure Code Review with Mixture of Prompts. arXiv:2510.12186
@@
- Pham et al. (2024) Van-Hau Pham, Do Thi Thu Hien, Hien Do Hoang, and Phan The Duy. 2024. Defect-Scanner: A Comparative Empirical Study on Language Model and Deep Learning Approach for Software Vulnerability Detection. International Journal of Information Security (IJIS) 23 (2024), 1–14. Issue 6. doi:10.1007/s10207-024-00901-4
@@
- Prasad et al. (2025) Rupesh Prasad, Lokesh B, Swathika Nivedini D, Piyush Pratap Singh, Arvind Maurya, and Atul Singh. 2025. Sink Vulnerability Type Prediction Using Small Language Model (SLM). In International Conference on Cognitive Computing in Engineering, Communications, Sciences and Biomedical Health Informatics (IC3ECSBHI). IEEE, Greater Noida, India, 496–501. doi:10.1109/IC3ECSBHI63591.2025.10991300
@@
- Qi et al. (2024) Weiliang Qi, Jiahao Cao, Darsh Poddar, Sophia Li, and Xinda Wang. 2024. Software Vulnerability Detection using Large Language Models. In International Conference on Security and Privacy in Communication Systems (SecureComm). Springer, Dubai, United Arab Emirates, 184–203. doi:10.1007/978-3-031-94458-1_9
- Qian et al. (2025) Zhuoyun Qian, Fangtian Zhong, Qin Hu, Yili Jiang, Jiaqi Huang, Mengfei Ren, and Jiguo Yu. 2025. Software Vulnerability Analysis Across Programming Language and Program Representation Landscapes: A Survey. arXiv:2503.20244
@@
- Ruohonen (2021) Jukka Ruohonen. 2021. The Similarities of Software Vulnerabilities for Interpreted Programming Languages. In IEEE International Conference on Progress in Informatics and Computing (PIC). IEEE, Shanghai, China, 304–307. doi:10.1109/PIC53636.2021.9687053
@@
- Safdar et al. (2025b) Rijha Safdar, Danyail Mateen, Syed Taha Ali, M. Umer Ashfaq, and Wajahat Hussain. 2025b. Data and Context Matter: Towards Generalizing AI-based Software Vulnerability Detection. arXiv:2508.16625
- Safdar et al. (2025a) Rijha Safdar, Danyail Mateen, Syed Taha Ali, and Wajahat Hussain. 2025a. Real-VulLLM: An LLM Based Assessment Framework in the Wild. arXiv:2510.04056
@@
- Sayagh and Ghafari (2025) Mohammed Sayagh and Mohammad Ghafari. 2025. Think Broad, Act Narrow: CWE Identification with Multi-Agent Large Language Models. arXiv:2508.01451
@@
- Shashwat et al. (2024) K Shashwat, F Hahn, X Ou, D Goldgof, J Ligatti, L Hall, R Rajagopalan, and A Tabari. 2024. A Preliminary Study on Using Large Language Models in Software Pentesting. In Network and Distributed System Security (NDSS) Symposium: Workshop on SOC Operations and Construction. NDSS Symposium, Marina del Rey, CA, USA.
- Sheng et al. (2025) Ze Sheng, Zhicheng Chen, Shuning Gu, Heqing Huang, Guofei Gu, and Jeff Huang. 2025. LLMs in Software Security: A Survey of Vulnerability Detection Techniques and Insights. Comput. Surveys 58, 5, Article 134 (2025), 35 pages. doi:10.1145/3769082
@@
- Simoni et al. (2025) Marco Simoni, Aleksandar Fontana, Giulio Rossolini, and Andrea Saracino. 2025. Improving LLM Reasoning for Vulnerability Detection via Group Relative Policy Optimization. arXiv:2507.03051
@@
- Sun et al. (2025c) Xiaobing Sun, Mingxuan Zhou, Sicong Cao, Xiaoxue Wu, Lili Bo, Di Wu, Bin Li, and Yang Xiang. 2025c. HgtJIT: Just-in-Time Vulnerability Detection Based on Heterogeneous Graph Transformer. IEEE Transactions on Dependable and Secure Computing (TDSC) 22, 6 (2025), 1–17. doi:10.1109/TDSC.2025.3586669
- Sun et al. (2025b) Yuqiang Sun, Daoyuan Wu, Yue Xue, Han Liu, Wei Ma, Lyuye Zhang, Yang Liu, and Yingjiu Li. 2025b. LLM4Vuln: A Unified Evaluation Framework for Decoupling and Enhancing LLMs’ Vulnerability Reasoning. arXiv:2401.16185
- Sun et al. (2025a) Zhihong Sun, Jia Li, Yao Wan, Chuanyi Li, Hongyu Zhang, Zhi jin, Ge Li, Hong Liu, Chen Lyu, and Songlin Hu. 2025a. Ensembling Large Language Models for Code Vulnerability Detection: An Empirical Evaluation. arXiv:2509.12629
@@
- Takieldeen et al. (2025) Ali E. Takieldeen, Merna Gamal, Abdelrahman Ahmed Salah, Mohamed Walid, Abdelrahman Mohamed Mahmoud, khaled Ahmed, Saleh Tamer Abdallah, Nada Samir Elsayed, and Aida A. Nasr. 2025. AI-Powered Vulnerability Detection in Code Using BERT-Based LLM with Transparency Measures. In International Telecommunications Conference (ITC-Egypt). IEEE, Cairo, Egypt, 297–304. doi:10.1109/ITC-Egypt66095.2025.11186618
@@
- Tian and Zhang (2025) Lei Tian and Cheng Zhang. 2025. EFVD: A Framework of Source Code Vulnerability Detection via Fusion of Enhanced Graph Representation Learning and Pre-trained Transformer-Based Model. In International Conference on Computer Network Security and Software Engineering (CNSSE). ACM, Qingdao, China, 316–320. doi:10.1145/3732365.3732421
@@
- Tsai et al. (2025a) Chung-Nan Tsai, Xin Wang, Cheng-Hsiung Lee, and Ching-Sheng Lin. 2025a. A Sequential Multi-Stage Approach for Code Vulnerability Detection via Confidence- and Collaboration-based Decision Making. In Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Suzhou, China, 21162–21168. doi:10.18653/v1/2025.emnlp-main.1071
- Tsai et al. (2025b) Chung-Nan Tsai, Jingnan Xie, Chun-Ming Lai, and Ching-Sheng Lin. 2025b. Leveraging Intra-and Inter-References in Vulnerability Detection using Multi-Agent Collaboration Based on LLMs. Cluster Computing 28, 15 (2025), 1–12. doi:10.1007/s10586-025-05721-2
@@
- Wang et al. (2024b) Rongcun Wang, Senlei Xu, Yuan Tian, Xingyu Ji, Xiaobing Sun, and Shujuang Jiang. 2024b. SCL-CVD: Supervised Contrastive Learning for Code Vulnerability Detection via GraphCodeBERT. Computers & Security (COSE) 145 (2024), 103994. doi:10.1016/j.cose.2024.103994
- Wang et al. (2025g) Weizhou Wang, Eric Liu, Xiangyu Guo, Xiao Hu, Ilya Grishchenko, and David Lie. 2025g. ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data. arXiv:2408.16028